# Slow is Fast! Dissecting Ethereum’s Slow Liquidity Drain Scams

### v2

@@ Abstract
We identify the slow liquidity drain (SLID) scam, an insidious and highly profitable threat to decentralized finance (DeFi), posing a large-scale, persistent, and growing risk to the ecosystem. Unlike traditional scams such as rug pulls or honeypots (USENIX Sec’19, USENIX Sec’23), SLID gradually siphons funds from liquidity pools over extended periods, making detection significantly more challenging. In this paper, we conducted the first large-scale empirical analysis of 319,166 liquidity pools across six major decentralized exchanges (DEXs) since 2018. We identified 3,117 SLID affected liquidity pools, resulting in cumulative losses of more than US\$103 million. We propose a rule-based heuristic and an enhanced machine learning model for early detection. Our machine learning model achieves a detection speed 4.77 times faster than the heuristic while maintaining 95% accuracy. Our study establishes a foundation for protecting DeFi investors at an early stage and promoting transparency in the DeFi ecosystem.
1 Introduction
@@
A pivotal moment for DeFi occurred with the rise of Uniswap [7], the leading decentralized exchange (DEX) on Ethereum. Uniswap introduced the Automated Market Maker (AMM) [8, 9, 10], enabling users to create liquidity pools (LPs) [^1] We use the abbreviation LP to denote liquidity pool, contrasting with sources [1, 11] that define it as liquidity provider. using their newly created tokens and facilitating direct asset exchanges without reliance on third parties like centralized exchanges (CEXs). Subsequent DEXs, such as Curve [12] and SushiSwap [13], further attracted investors by offering a broad range of cryptocurrencies, including SHIB [14] and PEPE [15]. The market value of DEXs [16, 17, 18] surged from \$587K in 2018 to \$7.71B during the “DeFi summer” of 2020 and eventually surpassed \$100B in 2021 [19]. To date, DEXs handle billions of dollars in daily trading volume [20].
@@
Decentralized exchanges have unsurprisingly become prime targets for financial scams [21, 22]. De.fi reported [23] a staggering \$1.39 billion loss in Ethereum’s DeFi sector due to scams. These scams often involve the rapid creation of liquidity pools, allowing fraudsters to retain full control over their value. Once users invest, they find themselves unable to sell or recover their funds. Two infamous examples include the rug pull [24], where developers abruptly abandon a project after withdrawing liquidity, and the honeypot [25], which lures users by permitting deposits but blocking withdrawals.
@@
Although advancements in scam detection and increased user awareness have reduced the prevalence of such straightforward scams, fraudsters are now resorting to more sophisticated schemes designed to bypass detection mechanisms and exploit unsuspecting users.
@@
Consider the case of a GoPlus report that flagged the Emily token as a potential rug pull due to the presence of an “unlimited mint function” [26]. However, upon closer examination, the token did not display the typical characteristics of a rug pull. This observation demanded further scrutiny, leading to the discovery of a far more insidious and pervasive threat within DeFi – a new and covert scam which we term the slow liquidity drain (SLID) scam. This phenomenon represents a sophisticated evolution in fraudulent tactics, exploiting the trust and complexity of decentralized finance to remain hidden. Unlike traditional DEX scams that exhibit clear characteristics (e.g., abrupt withdrawal, sale restriction), the SLID scam operates without disrupting the normal functionality of the liquidity pool. Instead, it gradually siphons assets through a slow but calculated profit-taking strategy. This stealthy approach enables the scam to blend with legitimate pools, making detection more challenging. As a result, investors may experience prolonged losses without immediate realization.
@@
Our analysis exposes the substantial fraudulent profits generated by SLID scams. Between 2018 and 2022, SLID scams siphoned off over \$64.33M in realized illicit gains, accounting for 14.3% of the total decentralized exchange (DEX) scam losses on Ethereum and Binance Chain combined (estimated at \$400M [27]). This striking figure underscores the broad impact of SLID scams within the DeFi ecosystem. Extending the scope to cover 2018 through 2024, these scams have inflicted a cumulative loss of \$103M, including realized and unrealized profits, marking them as a persistent threat to DeFi. Furthermore, our analysis of the dataset collected from Ethereum [28] reveals a yearly increase in the number of SLID scams since their emergence in 2018 (Figure 9). This trend highlights the urgency of raising awareness about SLID scams to mitigate the growing threat they pose to DeFi ecosystems.
@@
    RQ1. What are the common features of SLID scams? Currently, SLID scams lack formal recognition. Their defining characteristics remain unexplored. We aim to establish a clear definition of these scams, aiding in the identification of similar patterns in SLID liquidity pools and facilitating their distinction from legitimate ones.
@@
    RQ2. Can we build and evaluate a rule-based detector to identify SLID scams? Using the identified characteristics of SLID scams, we aim to design a rule-based heuristic to detect SLID liquidity pools on Ethereum. This heuristic will be evaluated in conjunction with external data from legitimate liquidity pools. Additionally, we aim to gain deeper insights into SLID scams by analyzing cases with unexpected outcomes (i.e., false positives, false negatives).
@@
    RQ3. Can we use the heuristic to automatically detect SLID scams at a large scale? Based on RQ2, we aim to apply the heuristic at scale to identify and label all SLID scams across the six largest Ethereum DEXs. However, the heuristic’s dependency on a long historical range of transaction data introduces latency in flagging SLID scams. To address this limitation, we also aim to develop a more proactive method capable of providing early and rapid warnings.
We first established our initial ground truth dataset by identifying community-reported rug pull cases that later failed to trigger the proposed rug pull detectors and exhibited characteristics similar to our research motivation. This process resulted in the identification of 71 liquidity pools that matched these descriptions. We then conducted a pilot study to analyze common scam behaviors shared by this initial ground truth dataset (§5). Based on the insights extracted from this analysis, we designed an efficient rule-based heuristic capable of identifying SLID liquidity pools by examining their DEX activity records and smart contract security features. This heuristic was tested using legitimate liquidity pools, with the evaluation outcomes discussed in detail (§6).
Figure 1: The workflow of our SLID scam: ① create a liquidity pool; ② add tokens (base & paired); ③ avoid burning initial shares/tokens; ④ marketing campaign; ⑤ invest funds and buy tokens; ⑥ normal activities (buy orders); ⑦ slowly withdraw/sell funds; ⑧ liquidity drop; ⑨ money loss.
Next, we applied the proposed rule-based heuristic to 319,166 Ethereum liquidity pools across the six largest DEXs on Ethereum (§7). This resulted in the flagging of 3,117 SLID scams, which collectively generated a profit of \$103.2M on these DEXs as of July 13, 2024. This analysis introduced the first-ever ground truth dataset of confirmed SLID scams. We then analyzed the flagged instances through three key dimensions: vitality, trend analysis, and profit, providing deeper insights into the tactics and the overall impact of SLID scams. Leveraging this expanded ground truth, we applied three ensemble machine learning (ML) algorithms (§7.3) of Random Forest, XGBoost, and Logistic Regression for early detection of SLID scams, capable of identifying potential scams with a shorter required DEX activity history that can detect SLID scam 4.77 times faster than the heuristic while maintaining high detection accuracy of 95%.
@@
    We applied our heuristic to six major DEXs on the Ethereum network, identifying 3,117 liquidity pools exhibiting SLID behaviors (with responsible disclosure). We produced the first ground truth dataset for SLID scams, publicly released at https://anonymous.4open.science/r/SLID-sourcecode-0DA3, and enabled analysis across vitality, profit, and trend metrics.
@@
    To enable early warning, we further applied ML algorithms using our ground truth dataset. We successfully reduced detection time by 4.77x while maintaining 95% accuracy.
@@ 2 Technical Warm-Ups
Ethereum tokens. Ethereum [28, 29] is the second leading blockchain platform after Bitcoin [30]. Ethereum introduced smart contracts and enabled the automatic execution of code. Users can create tradeable digital assets on-chain, known as tokens. These tokens serve multiple purposes, including fueling decentralized applications (DApps), acting as shares/stakes in projects, or representing ownership in organizations. Most tokens on Ethereum adhere to ERC20 [31], which defines a set of functions and interfaces. The standard ensures interoperability, allowing smooth token transfers among users.
@@
Ethereum accounts are tied to unique key pairs and are categorized into two types [32]: (i) Externally owned accounts (EOAs) store tokens and facilitate interactions with other accounts; (ii) contract accounts are associated with smart contracts to provide specialized on-chain services.
@@
DEXs and liquidity pool.  Uniswap [33], the unicorn DEX on Ethereum, is the vital DApp that has significantly contributed to the growth of the DeFi ecosystem. Their AMM infrastructure [9] enables users to list their newly created tokens for trading by deploying a liquidity pool. A liquidity pool is a smart contract on the DEX that holds two or more tokens, typically including at least one well-known token with independent value (such as ETH, USDT, or USDC), referred to as the base token. The other token is another base token or a newly introduced token, known as the paired token.
@@
When initializing a liquidity pool, the owner must deposit a specific amount of both tokens. The value of this LP is determined by the total USD value of the base token it holds, while the price of the paired token is positively correlated to the pool’s overall value, following the provided AMM equations [34]. Trading the base token for the paired token is termed a buy, while trading in the opposite direction is a sell. These swap actions impact the pool’s dynamics: a buy increases the price of the traded token as more base tokens flow in, whereas a sell reduces the price as tokens are withdrawn.
@@
Any user can contribute to the pool’s liquidity via deposit with a certain amount of both tokens to earn incentives later or through withdraw to remove their previously provided liquidity. All activities are publicly recorded on-chain and are referred to as DEX orders or DEX activities of that pool.
@@
Legitimate LPs conduct the action of burning (i.e., permanent locking) initial LP tokens after deployment. This process removes the pool owner’s ability to control the liquidity, preventing them from taking unethical actions with investor funds, such as draining the pool. Burning LP tokens is widely regarded as an industry standard for establishing trustworthiness and transparency in DeFi space [33, 1, 2, 35].
@@
Rug pull.  The most common DEX scam leveraging liquidity pools is rug pull [24]. The scammer creates (so that owns) the scam pool and extracts all the base tokens. There are two basic operations through which scammers extract profits:
@@
    Sell: first, sell a large inflated amount of paired tokens (they owned) in exchange for all pool’s base tokens;
@@
    Withdraw: withdraw the liquidity they initially provided, “pulling” out all of the base tokens from the pool.
@@
When scammers engage in sell or withdraw activities and the value of the assets they retrieve exceeds their initial investment in the scam pool, the excess amount is defined as the scammer’s realized profits. Correspondingly, unrealized profits represents the remaining value of assets within the scam pool that still belong to the scammer but have not yet been converted into realized profits via sell or withdraw actions.
Those sudden profit-taking actions caused a massive drop in the paired token price and the pool’s trading volume, serving as common signals for detecting rug pull.

### v3

@@ Abstract
We identify the slow liquidity drain (SLID) scam, an insidious and highly profitable threat to decentralized finance (DeFi), posing a large-scale, persistent, and growing risk to the ecosystem. Unlike traditional scams such as rug pulls or honeypots (USENIX Sec’19, USENIX Sec’23), SLID gradually siphons funds from liquidity pools over extended periods, making detection significantly more challenging. In this paper, we conduct the first systematic measurement study of SLID scams among 319,166 liquidity pools across six major decentralized exchanges (DEXs). We identified 3,117 SLID affected liquidity pools, resulting in cumulative losses of more than US\$103 million. We propose a rule-based heuristic and an enhanced machine learning model for early detection. Our machine learning model achieves a detection speed 4.77 times faster than the heuristic while maintaining the F1-Score of 0.93. Our study establishes a foundation for protecting DeFi investors at an early stage and promoting transparency in the DeFi ecosystem.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.[23||24]xxxx www.ndss-symposium.org
I Introduction
@@
A pivotal moment for DeFi occurred with the rise of Uniswap [7], the leading DEX on Ethereum. Uniswap introduced the Automated Market Maker (AMM) [8, 9, 10], enabling users to create liquidity pools (LPs) [^1] Within this paper, we use the abbreviation LP to denote the liquidity pool, contrasting with sources [1, 11] that define it as the liquidity provider. using their newly created tokens and facilitating direct asset exchanges without reliance on third parties like centralized exchanges (CEXs). Subsequent DEXs, such as Curve [12] and SushiSwap [13], further attracted investors by offering a broad range of cryptocurrencies, including SHIB [14] and PEPE [15]. The market value of DEXs [16, 17, 18] surged from \$587K in 2018 to \$7.71B during the “DeFi summer” of 2020 and eventually surpassed \$100B in 2021 [19]. To date, DEXs handle billions of dollars in daily trading volume [20].
@@
DEXs unsurprisingly became prime targets for financial scams [21, 22]. De.fi reported [23] a staggering \$1.39b loss in Ethereum’s DeFi sector due to scams. These scams often involve the rapid creation of LPs, allowing fraudsters to retain full control over monetary value. Once invested, users are unable to sell or recover funds. Two infamous examples include rug pull [24] (developers abruptly abandon a project after withdrawing liquidity), and honeypot [25] (luring users by permitting deposits but blocking withdrawals).
@@
Although increased user awareness reduced the prevalence of such straightforward scams, fraudsters are now resorting to more sophisticated schemes designed to bypass detection mechanisms and exploit unsuspecting users. Consider the case of a GoPlus report that flagged the Emily token as a potential exploitation due to the presence of an “unlimited mint function” [26]. However, upon closer examination, the token did not display typical characteristics that are close to any of the existing DeFi scam types.
@@
This observation demanded further scrutiny, leading to this paper discovering a far more pervasive threat which we term the slow liquidity drain (SLID) scam. Unlike traditional DEX scams that exhibit clear characteristics (e.g., abrupt withdrawal, sale restriction), SLID operates without disrupting the normal functionality of LPs. Instead, it gradually siphons assets through a slow but calculated profit-taking strategy. This stealthy approach enables the scam to blend with legitimate LPs, making detection more challenging. Investors may experience prolonged losses without immediate realization.
@@
Our analysis exposes the substantial fraudulent profits generated by SLID scams. Between 2018 and 2022, SLID scams siphoned off over \$64.33M in realized illicit gains, accounting for 14.3% of the total DEX scam losses on Ethereum and Binance Chain combined (estimated at \$400M [27]). This striking figure underscores the broad impact of SLID scams within the DeFi ecosystem. Extending the scope to cover 2018 through 2024, these scams have inflicted a cumulative loss of \$103M, including realized and unrealized profits, marking them as a persistent threat to DeFi. Furthermore, our analysis of the dataset collected from Ethereum [28] reveals a yearly increase in the number of SLID scams since their emergence in 2018 (Figure 12). This trend highlights the urgency of raising awareness about SLID scams to mitigate the growing threat they pose to DeFi ecosystems.
@@
    RQ1. What are the common features of SLID scams? SLID scams lack formal recognition. Their defining characteristics remain unexplored. We aim to establish a formal definition of these scams, aiding in the identification of similar patterns in SLID LPs and facilitating their distinction from legitimate ones.
@@
    RQ2. Can we build and evaluate a rule-based detector to identify SLID scams? Using the identified characteristics of SLID scams, we aim to design a rule-based heuristic to detect SLID LPs on Ethereum. This heuristic will be evaluated in conjunction with external data from legitimate LPs. Additionally, we aim to gain deeper insights into SLID scams by analyzing cases with unexpected outcomes (i.e., false positives, false negatives).
@@
    RQ3. Can we use the heuristic to automatically detect SLID scams at a large scale? Building on RQ2, we aim to deploy the heuristic at scale to identify and label SLID scams across the six largest Ethereum DEXs. However, the heuristic’s reliance on a long historical window introduces detection latency, and its inherently fuzzy design may fail to capture borderline cases. To address these limitations, we propose developing a proactive detection framework that enables earlier warning signals and improves coverage of ambiguous or borderline SLID instances.
We first established our initial SLID-like dataset by identifying community-reported DeFi scam cases that were uncategorized and exhibited characteristics similar to our research motivation. This process resulted in the identification of 71 LPs that matched these descriptions. We then conducted a pilot study to analyze common scam behaviors shared by this initial SLID-like dataset (§IV). Based on the insights extracted from this analysis, we designed an efficient rule-based heuristic capable of identifying SLID LPs by examining their DEX activity records and smart contract security features. This heuristic was validated using legitimate LPs, with evaluation outcomes discussed in detail (§V).
Next, we applied our rule-based heuristic to 319,166 Ethereum LPs across the six largest DEXs on Ethereum (§VI). This resulted in the flagging of 3,117 SLID scams, which collectively generated a profit of \$103.2M on these DEXs as of July 13, 2024. This analysis introduced the first-ever ground truth dataset of confirmed SLID scams. We then analyzed the flagged instances through three key dimensions: vitality, trend analysis, and profit, providing deeper insights into the tactics and the overall impact of SLID scams. Leveraging this discovery of SLID ground truth, we applied three machine learning (ML) algorithms (§VI-C) of Random Forest, XGBoost, and Logistic Regression for early detection of SLID scams, capable of detect the borderline cases that have been left from our SLID fuzzy heuristic and identifying potential scams with such a shorter required DEX activity history that can detect SLID scam 4.77 times faster than the heuristic while maintaining high detection F1-Score of 0.93.
@@
    We applied our heuristic to six major DEXs on the Ethereum network, identifying 3,117 LPs exhibiting SLID behaviors (with responsible disclosure). We produced the first ground truth dataset for SLID scams, publicly released at https://anonymous.4open.science/r/SLID-sourcecode-41AD, and enabled analysis across vitality, profit, and trend metrics. We also responsibly disclosed our findings to affected DEX teams and made all our coding resources used in this study publicly available [29].
@@
    To enable early warning and discover borderline cases outside of the heuristic scope, we further applied ML algorithms via our ground truth dataset. We successfully reduced detection time by 4.77x while maintaining a 0.93 F1-score.
@@ II Technical Warm-Ups
Ethereum tokens. Ethereum [28, 30] is the 2nd leading blockchain after Bitcoin [31]. Ethereum introduced smart contracts and enabled the automatic execution of code. Users can create tradeable digital assets on-chain, known as tokens. Those tokens are fungible (adhering to ERC20 [32]) and serve multiple purposes, including fuelling decentralised applications (DApps), acting as stakes, or representing ownership.
@@
Ethereum accounts are tied to unique key pairs in two types [33]: (i) externally owned accounts (EOAs) store tokens and facilitate interactions with other accounts; (ii) contract accounts are associated with smart contracts to provide specialized on-chain services.
@@
DEXs and liquidity pool.  Uniswap [34], the unicorn DEX on Ethereum, is the vital DApp that has significantly contributed to the growth of the DeFi ecosystem. Their AMM infrastructure [9] enables users to list their newly created tokens for trading by deploying a liquidity pool (LP). A LP is a smart contract on DEXs that holds two (or more) tokens, typically including at least one well-known token with independent value (e.g., ETH, USDT, or USDC), referred to as the base token. The other token is another base token or a newly introduced token, known as the paired token.
@@
When initializing a LP, the owner must deposit a specific amount of both tokens. The value of this LP is determined by the total USD value of the base token it holds, while the price of the paired token is positively correlated to the LP’s overall value, following the provided AMM equations [35, 36]. Trading the base token for the paired token is termed a buy, while trading in the opposite direction is a sell. These swaps impact the LP’s dynamics: buy increases the price of the traded token as more base tokens flow in, whereas sell reduces prices as tokens are withdrawn.
@@
Any user can contribute to the LP’s liquidity via deposit with a certain amount of both tokens to earn incentives later or through withdraw to remove their previously provided liquidity. All activities are publicly recorded on-chain and are referred to as DEX orders or DEX activities of that LP.
When it is proven that the price of the paired token is correlated with the liquidity of the base token in the LP, the owner can guarantee financial safety when pulling the scam and always exit at a higher price to get profit (since when the trader comes, the liquidity of the base token first increases and leads to the increment of the price of the paired token, see more details at Appendix A-B).
@@
Legitimate LPs conduct the action of burning (i.e., permanently locking) initial LP tokens after deployment. This process removes the LP owner’s ability to control the liquidity, preventing them from taking unethical actions, such as draining the LP. Burning LP tokens is widely regarded as an industry standard for establishing trustworthiness and transparency in the DeFi space [1, 2, 34, 37].
@@
Rug pull.  The most common DEX scam leveraging LPs is rug pull [24]. The scammer creates (so that owns) the scam LP and withdraws all the base tokens in the LP later. There are two basic operations through which scammers extract profits:
@@
    Sell: first, sell a large inflated amount of paired tokens (they owned) in exchange for all LP’s base tokens;
@@
    Withdraw: withdraw the liquidity they initially provided, “pulling” out all of the base tokens from the LP.
@@
When scammers engage in sell or withdraw activities and the value of the retrieved assets exceeds their initial investment, those excess amounts are defined as the scammer’s realized profits. Correspondingly, unrealized profit represents the remaining value of assets within the scam LP that still belong to the scammer but have not yet been converted into realized profits via sell or withdraw actions.
Those profit-taking actions caused a sharp drop in the token’s price and trading volume. Investors incur immediate financial losses. New investors lose interest in the scam LP. These patterns serve as signals for detecting rug pulls.